\begin{thebibliography}{56}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Adi et~al.(2016)Adi, Kermany, Belinkov, Lavi, and
  Goldberg}]{DBLP:journals/corr/AdiKBLG16}
Yossi Adi, Einat Kermany, Yonatan Belinkov, Ofer Lavi, and Yoav Goldberg. 2016.
\newblock \href {http://arxiv.org/abs/1608.04207} {Fine-grained analysis of
  sentence embeddings using auxiliary prediction tasks}.
\newblock \emph{CoRR}, abs/1608.04207.

\bibitem[{Bau et~al.(2019)Bau, Belinkov, Sajjad, Durrani, Dalvi, and
  Glass}]{bau2018identifying}
Anthony Bau, Yonatan Belinkov, Hassan Sajjad, Nadir Durrani, Fahim Dalvi, and
  James Glass. 2019.
\newblock \href {https://openreview.net/forum?id=H1z-PsR5KX} {Identifying and
  controlling important neurons in neural machine translation}.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Belinkov et~al.(2017{\natexlab{a}})Belinkov, Durrani, Dalvi, Sajjad,
  and Glass}]{belinkov:2017:acl}
Yonatan Belinkov, Nadir Durrani, Fahim Dalvi, Hassan Sajjad, and James Glass.
  2017{\natexlab{a}}.
\newblock \href
  {https://aclanthology.coli.uni-saarland.de/pdf/P/P17/P17-1080.pdf} {{What do
  Neural Machine Translation Models Learn about Morphology?}}
\newblock In \emph{Proceedings of the 55th Annual Meeting of the Association
  for Computational Linguistics (ACL)}, Vancouver. Association for
  Computational Linguistics.

\bibitem[{Belinkov et~al.(2020{\natexlab{a}})Belinkov, Durrani, Dalvi, Sajjad,
  and Glass}]{belinkov-etal-2020-linguistic}
Yonatan Belinkov, Nadir Durrani, Fahim Dalvi, Hassan Sajjad, and James Glass.
  2020{\natexlab{a}}.
\newblock \href {https://doi.org/10.1162/coli_a_00367} {On the linguistic
  representational power of neural machine translation models}.
\newblock \emph{Computational Linguistics}, 46(1):1--52.

\bibitem[{Belinkov et~al.(2020{\natexlab{b}})Belinkov, Gehrmann, and
  Pavlick}]{belinkov-etal-2020-interpretability}
Yonatan Belinkov, Sebastian Gehrmann, and Ellie Pavlick. 2020{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-tutorials.1}
  {Interpretability and analysis in neural {NLP}}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics: Tutorial Abstracts}, pages 1--5, Online.
  Association for Computational Linguistics.

\bibitem[{Belinkov and Glass(2019)}]{belinkov-glass-2019-analysis}
Yonatan Belinkov and James Glass. 2019.
\newblock \href {https://doi.org/10.1162/tacl_a_00254} {Analysis methods in
  neural language processing: A survey}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  7:49--72.

\bibitem[{Belinkov et~al.(2017{\natexlab{b}})Belinkov, M{\`a}rquez, Sajjad,
  Durrani, Dalvi, and Glass}]{belinkov-etal-2017-evaluating}
Yonatan Belinkov, Llu{\'\i}s M{\`a}rquez, Hassan Sajjad, Nadir Durrani, Fahim
  Dalvi, and James Glass. 2017{\natexlab{b}}.
\newblock \href {https://www.aclweb.org/anthology/I17-1001} {Evaluating layers
  of representation in neural machine translation on part-of-speech and
  semantic tagging tasks}.
\newblock In \emph{Proceedings of the Eighth International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 1--10, Taipei,
  Taiwan. Asian Federation of Natural Language Processing.

\bibitem[{Conneau et~al.(2018)Conneau, Kruszewski, Lample, Barrault, and
  Baroni}]{conneau-etal-2018-cram}
Alexis Conneau, German Kruszewski, Guillaume Lample, Lo{\"\i}c Barrault, and
  Marco Baroni. 2018.
\newblock \href {https://doi.org/10.18653/v1/P18-1198} {What you can cram into
  a single {\$}{\&}!{\#}* vector: Probing sentence embeddings for linguistic
  properties}.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 2126--2136,
  Melbourne, Australia. Association for Computational Linguistics.

\bibitem[{Dalvi et~al.(2019{\natexlab{a}})Dalvi, Durrani, Sajjad, Belinkov,
  Bau, and Glass}]{dalvi:2019:AAAI}
Fahim Dalvi, Nadir Durrani, Hassan Sajjad, Yonatan Belinkov, D.~Anthony Bau,
  and James Glass. 2019{\natexlab{a}}.
\newblock What is one grain of sand in the desert? analyzing individual neurons
  in deep nlp models.
\newblock In \emph{Proceedings of the Thirty-Third AAAI Conference on
  Artificial Intelligence (AAAI, Oral presentation)}.

\bibitem[{Dalvi et~al.(2017)Dalvi, Durrani, Sajjad, Belinkov, and
  Vogel}]{dalvi:2017:ijcnlp}
Fahim Dalvi, Nadir Durrani, Hassan Sajjad, Yonatan Belinkov, and Stephan Vogel.
  2017.
\newblock {Understanding and Improving Morphological Learning in the Neural
  Machine Translation Decoder}.
\newblock In \emph{Proceedings of the 8th International Joint Conference on
  Natural Language Processing (IJCNLP)}.

\bibitem[{Dalvi et~al.(2019{\natexlab{b}})Dalvi, Nortonsmith, Bau, Belinkov,
  Sajjad, Durrani, and Glass}]{neurox-aaai19:demo}
Fahim Dalvi, Avery Nortonsmith, D.~Anthony Bau, Yonatan Belinkov, Hassan
  Sajjad, Nadir Durrani, and James Glass. 2019{\natexlab{b}}.
\newblock Neurox: A toolkit for analyzing individual neurons in neural
  networks.
\newblock In \emph{AAAI Conference on Artificial Intelligence (AAAI)}.

\bibitem[{Dalvi et~al.(2020)Dalvi, Sajjad, Durrani, and
  Belinkov}]{dalvi-2020-CCFS}
Fahim Dalvi, Hassan Sajjad, Nadir Durrani, and Yonatan Belinkov. 2020.
\newblock Analyzing redundancy in pretrained transformer models.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP-2020)}, Online. Association for
  Computational Linguistics.

\bibitem[{Danilevsky et~al.(2020)Danilevsky, Qian, Aharonov, Katsis, Kawas, and
  Sen}]{danilevsky-etal-2020-survey}
Marina Danilevsky, Kun Qian, Ranit Aharonov, Yannis Katsis, Ban Kawas, and
  Prithviraj Sen. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.aacl-main.46} {A survey
  of the state of explainable {AI} for natural language processing}.
\newblock In \emph{Proceedings of the 1st Conference of the Asia-Pacific
  Chapter of the Association for Computational Linguistics and the 10th
  International Joint Conference on Natural Language Processing}, pages
  447--459, Suzhou, China. Association for Computational Linguistics.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{devlin-etal-2019-bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock {BERT}: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, Minneapolis, Minnesota.
  Association for Computational Linguistics.

\bibitem[{Dhamdhere et~al.(2020)Dhamdhere, Agarwal, and
  Sundararajan}]{dhamdhere2020shapley}
Kedar Dhamdhere, Ashish Agarwal, and Mukund Sundararajan. 2020.
\newblock \href {http://arxiv.org/abs/1902.05622} {The shapley taylor
  interaction index}.

\bibitem[{Dhamdhere et~al.(2018{\natexlab{a}})Dhamdhere, Sundararajan, and
  Yan}]{DBLP:journals/corr/abs-1805-12233}
Kedar Dhamdhere, Mukund Sundararajan, and Qiqi Yan. 2018{\natexlab{a}}.
\newblock \href {http://arxiv.org/abs/1805.12233} {How important is a neuron?}
\newblock \emph{CoRR}, abs/1805.12233.

\bibitem[{Dhamdhere et~al.(2018{\natexlab{b}})Dhamdhere, Sundararajan, and
  Yan}]{cond}
Kedar Dhamdhere, Mukund Sundararajan, and Qiqi Yan. 2018{\natexlab{b}}.
\newblock \href {http://arxiv.org/abs/1805.12233} {How important is a neuron?}

\bibitem[{Durrani et~al.(2019)Durrani, Dalvi, Sajjad, Belinkov, and
  Nakov}]{durrani-etal-2019-one}
Nadir Durrani, Fahim Dalvi, Hassan Sajjad, Yonatan Belinkov, and Preslav Nakov.
  2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1154} {One size does not fit
  all: Comparing {NMT} representations of different granularities}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 1504--1516,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Durrani et~al.(2021)Durrani, Sajjad, and Dalvi}]{durrani-FT}
Nadir Durrani, Hassan Sajjad, and Fahim Dalvi. 2021.
\newblock How transfer learning impacts linguistic knowledge in deep nlp
  models?
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL 2021}, Online. Association for Computational Linguistics.

\bibitem[{Durrani et~al.(2020)Durrani, Sajjad, Dalvi, and
  Belinkov}]{durrani-etal-2020-analyzing}
Nadir Durrani, Hassan Sajjad, Fahim Dalvi, and Yonatan Belinkov. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.395} {Analyzing
  individual neurons in pre-trained language models}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 4865--4880, Online. Association
  for Computational Linguistics.

\bibitem[{Gu et~al.(2021)Gu, Feng, and Xie}]{gu2021pruningthenexpanding}
Shuhao Gu, Yang Feng, and Wanying Xie. 2021.
\newblock \href {http://arxiv.org/abs/2103.13678} {Pruning-then-expanding model
  for domain adaptation of neural machine translation}.

\bibitem[{Gulordava et~al.(2018)Gulordava, Bojanowski, Grave, Linzen, and
  Baroni}]{gulordava-etal-2018-colorless}
Kristina Gulordava, Piotr Bojanowski, Edouard Grave, Tal Linzen, and Marco
  Baroni. 2018.
\newblock \href {https://doi.org/10.18653/v1/N18-1108} {Colorless green
  recurrent networks dream hierarchically}.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, pages 1195--1205, New Orleans,
  Louisiana. Association for Computational Linguistics.

\bibitem[{Hewitt and Liang(2019)}]{hewitt-liang-2019-designing}
John Hewitt and Percy Liang. 2019.
\newblock \href {https://doi.org/10.18653/v1/D19-1275} {Designing and
  interpreting probes with control tasks}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 2733--2743, Hong Kong,
  China. Association for Computational Linguistics.

\bibitem[{Janizek et~al.(2020)Janizek, Sturmfels, and
  Lee}]{janizek2020explaining}
Joseph~D. Janizek, Pascal Sturmfels, and Su-In Lee. 2020.
\newblock \href {http://arxiv.org/abs/2002.04138} {Explaining explanations:
  Axiomatic feature interactions for deep networks}.

\bibitem[{Jumelet and Hupkes(2019)}]{diagnnose}
Jaap Jumelet and Dieuwke Hupkes. 2019.
\newblock \href {https://doi.org/10.5281/zenodo.3445477} {diagnnose: A neural
  net analysis library}.

\bibitem[{K{\'a}d{\'a}r et~al.(2017)K{\'a}d{\'a}r, Chrupa{\l}a, and
  Alishahi}]{kadar-etal-2017-representation}
{\'A}kos K{\'a}d{\'a}r, Grzegorz Chrupa{\l}a, and Afra Alishahi. 2017.
\newblock \href {https://doi.org/10.1162/COLI_a_00300} {Representation of
  linguistic form and function in recurrent neural networks}.
\newblock \emph{Computational Linguistics}, 43(4):761--780.

\bibitem[{Karpathy et~al.(2015)Karpathy, Johnson, and
  Fei-Fei}]{karpathy2015visualizing}
Andrej Karpathy, Justin Johnson, and Li~Fei-Fei. 2015.
\newblock \href {http://arxiv.org/abs/1506.02078} {Visualizing and
  understanding recurrent networks}.

\bibitem[{Kokhlikyan et~al.(2020)Kokhlikyan, Miglani, Martin, Wang, Alsallakh,
  Reynolds, Melnikov, Kliushkina, Araya, Yan, and
  Reblitz-Richardson}]{kokhlikyan2020captum}
Narine Kokhlikyan, Vivek Miglani, Miguel Martin, Edward Wang, Bilal Alsallakh,
  Jonathan Reynolds, Alexander Melnikov, Natalia Kliushkina, Carlos Araya, Siqi
  Yan, and Orion Reblitz-Richardson. 2020.
\newblock \href {http://arxiv.org/abs/2009.07896} {{Captum: A unified and
  generic model interpretability library for PyTorch}}.

\bibitem[{Lakretz et~al.(2019)Lakretz, Kruszewski, Desbordes, Hupkes, Dehaene,
  and Baroni}]{lakretz-etal-2019-emergence}
Yair Lakretz, German Kruszewski, Theo Desbordes, Dieuwke Hupkes, Stanislas
  Dehaene, and Marco Baroni. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1002} {The emergence of number
  and syntax units in {LSTM} language models}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 11--20, Minneapolis,
  Minnesota. Association for Computational Linguistics.

\bibitem[{Li et~al.(2016{\natexlab{a}})Li, Chen, Hovy, and
  Jurafsky}]{li-etal-2016-visualizing}
Jiwei Li, Xinlei Chen, Eduard Hovy, and Dan Jurafsky. 2016{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/N16-1082} {Visualizing and
  understanding neural models in {NLP}}.
\newblock In \emph{Proceedings of the 2016 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 681--691, San Diego, California. Association for
  Computational Linguistics.

\bibitem[{Li et~al.(2016{\natexlab{b}})Li, Monroe, and Jurafsky}]{LiMJ16a}
Jiwei Li, Will Monroe, and Dan Jurafsky. 2016{\natexlab{b}}.
\newblock \href {http://arxiv.org/abs/1612.08220} {Understanding neural
  networks through representation erasure}.
\newblock \emph{CoRR}, abs/1612.08220.

\bibitem[{Lillian et~al.(2018)Lillian, Meyes, and
  Meisen}]{DBLP:journals/corr/abs-1812-05687}
Peter Lillian, Richard Meyes, and Tobias Meisen. 2018.
\newblock \href {http://arxiv.org/abs/1812.05687} {Ablation of a robot's brain:
  Neural networks under a knife}.
\newblock \emph{CoRR}, abs/1812.05687.

\bibitem[{Linzen et~al.(2016)Linzen, Dupoux, and Goldberg}]{linzen_tacl}
Tal Linzen, Emmanuel Dupoux, and Yoav Goldberg. 2016.
\newblock {Assessing the ability of LSTMs to learn syntax-sensitive
  dependencies}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  4:521– 535.

\bibitem[{Lipton(2016)}]{lipton2016mythos}
Zachary~C Lipton. 2016.
\newblock {The Mythos of Model Interpretability}.
\newblock In \emph{ICML Workshop on Human Interpretability in Machine Learning
  (WHI)}.

\bibitem[{Lundberg and Lee(2017)}]{shappely_NIPS2017_7062}
Scott~M Lundberg and Su-In Lee. 2017.
\newblock \href
  {http://papers.nips.cc/paper/7062-a-unified-approach-to-interpreting-model-predictions.pdf}
  {A unified approach to interpreting model predictions}.
\newblock In I.~Guyon, U.~V. Luxburg, S.~Bengio, H.~Wallach, R.~Fergus,
  S.~Vishwanathan, and R.~Garnett, editors, \emph{Advances in Neural
  Information Processing Systems 30}, pages 4765--4774. Curran Associates, Inc.

\bibitem[{Marvin and Linzen(2018)}]{marvin-linzen-2018-targeted}
Rebecca Marvin and Tal Linzen. 2018.
\newblock \href {https://doi.org/10.18653/v1/D18-1151} {Targeted syntactic
  evaluation of language models}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 1192--1202, Brussels, Belgium.
  Association for Computational Linguistics.

\bibitem[{Michael et~al.(2020)Michael, Botha, and
  Tenney}]{michael-etal-2020-asking}
Julian Michael, Jan~A. Botha, and Ian Tenney. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.552} {Asking
  without telling: Exploring latent ontologies in contextual representations}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 6792--6812, Online. Association
  for Computational Linguistics.

\bibitem[{Mu and Andreas(2020)}]{Mu-Nips}
Jesse Mu and Jacob Andreas. 2020.
\newblock \href {http://arxiv.org/abs/2006.14032} {Compositional explanations
  of neurons}.
\newblock \emph{CoRR}, abs/2006.14032.

\bibitem[{Na et~al.(2019)Na, Choe, Lee, and Kim}]{Na-ICLR}
Seil Na, Yo~Joong Choe, Dong{-}Hyun Lee, and Gunhee Kim. 2019.
\newblock \href {http://arxiv.org/abs/1902.07249} {Discovery of natural
  language concepts in individual units of cnns}.
\newblock \emph{CoRR}, abs/1902.07249.

\bibitem[{Peters et~al.(2018)Peters, Neumann, Iyyer, Gardner, Clark, Lee, and
  Zettlemoyer}]{peters-etal-2018-deep}
Matthew Peters, Mark Neumann, Mohit Iyyer, Matt Gardner, Christopher Clark,
  Kenton Lee, and Luke Zettlemoyer. 2018.
\newblock Deep contextualized word representations.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, New Orleans, Louisiana. Association
  for Computational Linguistics.

\bibitem[{Poerner et~al.(2018)Poerner, Roth, and
  Sch{\"u}tze}]{poerner-etal-2018-interpretable}
Nina Poerner, Benjamin Roth, and Hinrich Sch{\"u}tze. 2018.
\newblock \href {https://doi.org/10.18653/v1/W18-5437} {Interpretable textual
  neuron representations for {NLP}}.
\newblock In \emph{Proceedings of the 2018 {EMNLP} Workshop {B}lackbox{NLP}:
  Analyzing and Interpreting Neural Networks for {NLP}}, pages 325--327,
  Brussels, Belgium. Association for Computational Linguistics.

\bibitem[{Rethmeier et~al.(2020)Rethmeier, Saxena, and
  Augenstein}]{rethmeier2019txray}
Nils Rethmeier, Vageesh~Kumar Saxena, and Isabelle Augenstein. 2020.
\newblock \href {http://www.auai.org/uai2020/proceedings/197\_main\_paper.pdf}
  {Tx-ray: Quantifying and explaining model-knowledge transfer in
  (un-)supervised {NLP}}.
\newblock In \emph{Proceedings of the Thirty-Sixth Conference on Uncertainty in
  Artificial Intelligence, {UAI} 2020, virtual online, August 3-6, 2020}, page
  197. {AUAI} Press.

\bibitem[{Ribeiro et~al.(2016)Ribeiro, Singh, and Guestrin}]{lime}
Marco~T{\'{u}}lio Ribeiro, Sameer Singh, and Carlos Guestrin. 2016.
\newblock \href {http://arxiv.org/abs/1602.04938} {"why should {I} trust you?":
  Explaining the predictions of any classifier}.
\newblock \emph{CoRR}, abs/1602.04938.

\bibitem[{Simonyan et~al.(2014)Simonyan, Vedaldi, and Zisserman}]{saliency}
Karen Simonyan, Andrea Vedaldi, and Andrew Zisserman. 2014.
\newblock \href {http://arxiv.org/abs/1312.6034} {Deep inside convolutional
  networks: Visualising image classification models and saliency maps}.

\bibitem[{Suau et~al.(2020)Suau, Zappella, and Apostoloff}]{suau2020finding}
Xavier Suau, Luca Zappella, and Nicholas Apostoloff. 2020.
\newblock \href {http://arxiv.org/abs/2005.07647} {Finding experts in
  transformer models}.
\newblock \emph{CoRR}, abs/2005.07647.

\bibitem[{Sundararajan et~al.(2017)Sundararajan, Taly, and Yan}]{ig}
Mukund Sundararajan, Ankur Taly, and Qiqi Yan. 2017.
\newblock \href {http://arxiv.org/abs/1703.01365} {Axiomatic attribution for
  deep networks}.

\bibitem[{Tenney et~al.(2019)Tenney, Das, and Pavlick}]{tenney-etal-2019-bert}
Ian Tenney, Dipanjan Das, and Ellie Pavlick. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1452} {{BERT} rediscovers the
  classical {NLP} pipeline}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 4593--4601, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Torroba~Hennigen et~al.(2020)Torroba~Hennigen, Williams, and
  Cotterell}]{torroba-hennigen-etal-2020-intrinsic}
Lucas Torroba~Hennigen, Adina Williams, and Ryan Cotterell. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.15} {Intrinsic
  probing through dimension selection}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 197--216, Online. Association for
  Computational Linguistics.

\bibitem[{Tran et~al.(2018)Tran, Bisazza, and Monz}]{tran2018importance}
Ke~Tran, Arianna Bisazza, and Christof Monz. 2018.
\newblock {The Importance of Being Recurrent for Modeling Hierarchical
  Structure}.
\newblock \emph{arXiv preprint arXiv:1803.03585}.

\bibitem[{Tsang et~al.(2020)Tsang, Rambhatla, and Liu}]{achipelago}
Michael Tsang, Sirisha Rambhatla, and Yan Liu. 2020.
\newblock \href {http://arxiv.org/abs/2006.10965} {How does this interaction
  affect me? interpretable attribution for feature interactions}.

\bibitem[{Valipour et~al.(2019)Valipour, Lee, Jamacaro, and
  Bessega}]{valipur-2019}
Mehrdad Valipour, En{-}Shiun~Annie Lee, Jaime~R. Jamacaro, and Carolina
  Bessega. 2019.
\newblock \href {http://arxiv.org/abs/1912.05308} {Unsupervised transfer
  learning via {BERT} neuron selection}.
\newblock \emph{CoRR}, abs/1912.05308.

\bibitem[{Vig et~al.(2020)Vig, Gehrmann, Belinkov, Qian, Nevo, Singer, and
  Shieber}]{vig_causal_genderbias_2020}
Jesse Vig, Sebastian Gehrmann, Yonatan Belinkov, Sharon Qian, Daniel Nevo,
  Yaron Singer, and Stuart Shieber. 2020.
\newblock {Causal Mediation Analysis for Interpreting Neural NLP: The Case of
  Gender Bias}.
\newblock \emph{arXiv preprint arXiv:2004.12265}.

\bibitem[{Voita et~al.(2020)Voita, Sennrich, and Titov}]{voita2020analyzing}
Elena Voita, Rico Sennrich, and Ivan Titov. 2020.
\newblock \href {http://arxiv.org/abs/2010.10907} {Analyzing the source and
  target contributions to predictions in neural machine translation}.

\bibitem[{Wallace et~al.(2020)Wallace, Gardner, and
  Singh}]{wallace-etal-2020-interpreting}
Eric Wallace, Matt Gardner, and Sameer Singh. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-tutorials.3}
  {Interpreting predictions of {NLP} models}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing: Tutorial Abstracts}, pages 20--23, Online.
  Association for Computational Linguistics.

\bibitem[{Wolf et~al.(2019)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac,
  Rault, Louf, Funtowicz, and Brew}]{Wolf2019HuggingFacesTS}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue,
  Anthony Moi, Pierric Cistac, Tim Rault, R'emi Louf, Morgan Funtowicz, and
  Jamie Brew. 2019.
\newblock Huggingface's transformers: State-of-the-art natural language
  processing.
\newblock \emph{ArXiv}, abs/1910.03771.

\bibitem[{Wu et~al.(2020)Wu, Belinkov, Durrani, Dalvi, and Glass}]{wu:2020:acl}
John Wu, Hassan Belinkov, Yonatan~Sajjad, Nadir Durrani, Fahim Dalvi, and James
  Glass. 2020.
\newblock {Similarity Analysis of Contextual Word Representation Models}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics (ACL)}, Seattle. Association for Computational
  Linguistics.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[]{naacl2021}

\usepackage{times}
\usepackage{latexsym}
\usepackage{enumerate}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\title{Fine-grained Interpretation and Causation Analysis in Deep NLP Models}

\author{
Hassan Sajjad ~~~ Narine Kokhlikyan\textsuperscript{*} ~~~ Fahim Dalvi ~~~ Nadir Durrani \\~\\
{\tt \{hsajjad,faimaduddin,ndurrani\}@hbku.edu.qa} \\ 
Qatar Computing Research Institute, HBKU Research Complex, Doha 5825, Qatar \\
{\tt narine@fb.com} \\
\textsuperscript{*}Facebook AI, 1 Facebook Way, Menlo Park, CA 94025, USA \\ 

}

\begin{document}
\maketitle
\section{Introduction}
Deep neural networks have constantly pushed the state-of-the-art performance in natural language processing (NLP) and are now considered as the de facto modeling approach in solving most complex NLP tasks such as machine translation, summarization and question-answering. Despite the benefits and the usefulness of deep neural networks at-large, their opaqueness is a major cause of concern. 
Interpreting neural networks is considered important for increasing trust in AI systems, providing additional information to decision makers, and assisting ethical decision making~\cite{lipton2016mythos}.

Work done on neural interpretation can be broadly divided into two types: i) intrinsic analysis, ii) causation analysis. The former class 
analyze representations learned within neural networks 
~\cite{DBLP:journals/corr/AdiKBLG16,belinkov:2017:acl,belinkov-etal-2017-evaluating,dalvi:2017:ijcnlp,conneau-etal-2018-cram,tenney-etal-2019-bert,durrani-etal-2019-one}.
The latter 
tries to understand the network with various behavioural studies 
~\cite{gulordava-etal-2018-colorless,linzen_tacl,marvin-linzen-2018-targeted}, 
or by studying importance of input features and neurons with respect to a  prediction~\cite{DBLP:journals/corr/abs-1805-12233,shappely_NIPS2017_7062,tran2018importance}. These works on interpretation have been presented in earlier tutorials, for example ACL 2020 \cite{belinkov-etal-2020-interpretability} and EMNLP 2020 \cite{wallace-etal-2020-interpreting}. The former focused on representation analysis and behavioral studies and the latter discusses work done on the causal interpertation such as assessing model's behavior using constructed examples. Both these great tutorials serve as a starting point for the new researchers in this area.


The representation analysis, also called as structural analysis, is useful to understand how various core linguistic properties are learned in the model \cite{belinkov-etal-2020-linguistic}. However, the analysis suffers from a few limitations. It  mainly focuses at interpreting full vector representations and does not study the role of fine-grained components in the representation i.e. neurons. Also the findings of representation analysis do not link with the cause of a prediction~\cite{belinkov-glass-2019-analysis}. While the behavioral analysis evaluates model predictions, it does not typically connect them with the influence of the input features and the internal components of a model~\cite{vig_causal_genderbias_2020}. 

In this tutorial, we 
present and discuss the research work on interpreting 
fine-grained components of a model 
%
%
from two perspectives, i) fine-grained interpretation, ii) causation analysis. The former 
introduces methods to analyze individual neurons and a group of neurons with respect to a 
language property or a task. The latter 
studies the role of neurons and input features in explaining decisions made by the model.
We will cover important research questions such as i) how is knowledge distributed across the model components? ii) what knowledge learned within the model is used for specific predictions? iii) does the inhibition of specific knowledge in the model change predictions? iv) how do different modeling and optimization choices impact the underlying knowledge?

Recent work on interpreting neurons has shown that in-addition to gaining better understanding of the inner workings of neural networks, the neuron-level interpretation has applications in model distillation \cite{rethmeier2019txray}, domain adaptation \cite{gu2021pruningthenexpanding} or efficient feature selection \cite{dalvi-2020-CCFS} e.g., by removing unimportant neurons, facilitating architecture search, and mitigating model bias by identifying neurons responsible for sensitive attributes like gender, race or politeness \cite{bau2018identifying,vig_causal_genderbias_2020}. 
These recent works are not only enabling better understanding of these networks, but are also leading towards better, fairer and more environmental-friendly models, which are all important goals for the 
Artificial Intelligence community at large. 

\section{Description}
\label{sec:description}

The tutorial 
is divided into two main parts: i) fine-grained interpretation, and ii) causation analysis. 
%
%
The first part of the tutorial
covers  methods that align neurons to human interpretable concepts or study the most salient neurons in the network. We cluster these methods into four groups i) Visualization Methods \cite{karpathy2015visualizing,li-etal-2016-visualizing}, ii) Corpus Selection \cite{kadar-etal-2017-representation,poerner-etal-2018-interpretable,Na-ICLR, Mu-Nips}, iii) Neuron Probing \cite{dalvi:2019:AAAI,lakretz-etal-2019-emergence,valipur-2019,durrani-etal-2020-analyzing,durrani-FT} and iv) Unsupervised Methods \cite{bau2018identifying,torroba-hennigen-etal-2020-intrinsic,wu:2020:acl,michael-etal-2020-asking}. 
We 
also discuss evaluation methods 
that are used to 
measure the effectiveness of an interpretation method, such as accuracy, control tasks~\cite{hewitt-liang-2019-designing} and ablation studies~\cite{LiMJ16a,DBLP:journals/corr/abs-1812-05687,dalvi:2019:AAAI,lakretz-etal-2019-emergence}. 
Moreover, we 
cover various applications of these methods that go beyond interpretation such as efficient transfer learning~\cite{dalvi-2020-CCFS}, controlling system's behavior \cite{bau2018identifying, suau2020finding}, generating explanations ~\cite{Mu-Nips} and domain adaptation \cite{gu2021pruningthenexpanding}.

The second part, \textit{Causation Analysis}, 
focuses on the methods that 
characterize the role of neurons and layers towards a specific prediction. More concretely, we 
discuss gradient and perturbation-based attribution algorithms such as Integrated Gradients~\cite{ig}, Layer Conductance~\cite{cond}, Saliency~\cite{saliency}, SHapley Additive exPlanations(SHAP)~\cite{shappely_NIPS2017_7062} and showcase how they can help us to identify important neurons in different layers of a deep neural network.
Besides that we 
also dive deep into more recent and advanced attribution algorithms that take feature or neuron interactions into account. More specifically, we 
look into Integrated Hessians~\cite{janizek2020explaining}, Shapely Taylor index~\cite{dhamdhere2020shapley} and Archipelago~\cite{achipelago}.


Lastly, we 
present various open source toolkits and libraries that provide implementation of notable techniques in the area. A few examples  
include: Captum~\cite{kokhlikyan2020captum}, InterpretML\footnote{https://interpret.ml/}, NeuroX~\cite{neurox-aaai19:demo}, Ecco\footnote{https://www.eccox.io/} and Diagnnose~\cite{diagnnose}. We 
give a walk-through of how some of these tools can be used for fine-grained interpretation and causation analysis. 
Throughout the tutorial, 
we critically evaluate 
the strengths and weakness of 
the presented methods, 
and 
discuss future directions. 

%

\section{Outline}

\begin{enumerate}[I.]
    \item \textbf{Introduction:} We 
    introduce the topic and motivate it by providing the vision of model interpretability, and how it leads towards fair and ethical models 
    and towards better generalization. We 
    then describe various forms of interpretation and 
    outline the scope of the tutorial (15 minutes).
    
    
    \item{\textbf{Fine-grained Interpretation:}} We 
    present and discuss the work on neuron-level interpretation. 
    (90 minutes)
    
    \begin{itemize}
    \item{Methods} (30 minutes)
    \item{Evaluation} (15 minutes)
    \item{Findings} (30 minutes)
    \item{Practical} (15 minutes)
    \end{itemize}
    
    \item{\textbf{Causation Analysis:}} In causation analysis we 
    present various methods on interpreting model predictions with respect to input features and individual neurons. (60 minutes)
    
    \begin{itemize}
    \item{Methods} (30 minutes)
    \item{Evaluation} (10 minutes)
    \item{Practical} (20 minutes)
    \end{itemize}
    
    
    
    
    \item \textbf{Concept-based Interpretation of Prediction:} This part 
    aims to 
    bridge the gap between fine-grained interpretation and causation analysis. We 
    discuss how fine-grained interpretation and causation analysis can be combined to establish concept-based interpretation of model predictions. 
    (10 minutes)
    
    \item \textbf{Discussion:} The last part 
    discusses the overall challenges that the current work faces and suggests future directions. (10 minutes)

\end{enumerate}

\section{Prerequisites}
We assume a basic knowledge of the deep learning and familiarity with the LSTM-based and transformer-based pre-trained models such as ELMo~\cite{peters-etal-2018-deep} and BERT~\cite{devlin-etal-2019-bert}. Additionally, some familiarity with natural language processing tasks such as, named entity tagging, natural language inference, etc. would be 
useful but not mandatory. We do not expect participants to have familiarity with the research on the interpretation and analysis of deep models. 
Familiarity with Python, Pytorch and Transformers library~\cite{Wolf2019HuggingFacesTS} would be 
useful to understand the practical part.

\section{Reading List}
\begin{itemize}
    \item In order to get an overview of the interpretation field, trainees may look at the following survey papers: \newcite{belinkov-glass-2019-analysis} and \newcite{danilevsky-etal-2020-survey}.
    \item Fine-grained analysis and its Applications: \newcite{,bau2018identifying,dalvi:2019:AAAI,Mu-Nips,suau2020finding} etc.
    \item Causation analysis: \newcite{shappely_NIPS2017_7062} provides an overview of various methods introduced in literature. For more details, see the following papers: ~\newcite{voita2020analyzing,ig,cond,lime,janizek2020explaining}
\end{itemize}
In addition to the above list, interested trainees may look at the papers mentioned in Section \ref{sec:description}.

\section{Instructor Information (Alphabetic order}

\noindent\textbf{Fahim Dalvi}, Software Engineer, Qatar Computing Research Institute, Qatar

\noindent Email: faimaduddin@hbku.edu.qa \\
\noindent Website: \url{https://fdalvi.github.io}
\\~\\
Fahim Dalvi is an experienced Software Engineer with a demonstrated history of working in the research industry and is currently employed at the Qatar Computing Research Institute. Fahim's research is centered around the intersection of Natural Language Processing and Deep Learning, and he has worked on wide variety of problems in these fields including Machine Translation, Language Modelling and Explainability in Deep Neural Networks. He also spends his time converting research into practical applications, with a focus on scalable web applications. Fahim also spends some time every year mentoring and teaching Deep Learning at Fall and Summer schools.
\\~\\
\noindent\textbf{Hassan Sajjad}, Senior Research Scientist, Qatar Computing Research Institute, Qatar

\noindent Email: hsajjad@hbku.edu.qa \\
\noindent Website: \url{https://hsajjad.github.io}
\\~\\
Hassan Sajjad is a Senior Research Scientist at the Qatar Computing Research Institute (QCRI), HBKU. His research interests include the interpretation of deep neural models, machine translation, domain adaptation, and natural language processing involving low-resource and morphologically-rich languages. His research work has been published in several prestigious venues such as CL, CSL, ICLR, ACL, NAACL and EMNLP. His work in collaboration with MIT and Harvard on the interpretation of deep models has also been featured in several tech blogs including MIT News. Hassan co-organized BlackboxNLP 2020, and the WMT 2019/2020 machine translation robustness task. He served as an area chair for the analysis and interpretability, NLP Application, and machine translation tracks at various *CL conferences. In addition, Hassan has been regularly teaching courses on deep learning internationally at various spring and summer schools. 
\\~\\
\noindent\textbf{Narine Kokhlikyan}, Research Scientist, Facebook AI

\noindent Email: narine@fb.com \\
\noindent Website: \url{https://www.linkedin.com/in/narine-k-88916721/}
\\~\\
Narine is a Research Scientist focusing on Model Interpretability as part of PyTorch team at Facebook. Her research interests include the understanding of Deep Neural Network internals and their predictions across different applications such as Natural Language Processing, Computer Vision and Recommender Systems.  In the recent years she gave talks and presented tutorials  on  Model Interpretability  at KDD 2020 and NeurIPS 2019.
Before joining Facebook Narine worked on Natural Language Processing, Time Series Analysis and numerical optimizations.
\\~\\
\noindent\textbf{Nadir Durrani}, Senior Research Scientist, Qatar Computing Research Institute, Qatar \\
\noindent Email: ndurrani@hbku.edu.qa \\
\noindent Website: \url{http://alt.qcri.org/~ndurrani/}
\\~\\
Nadir Durrani is a Research Scientist at the Arabic Language Technologies group at Qatar Computing Research Institute. His research interests include interpretation of neural networks, neural and statistical machine translation (with focus on reordering, domain adaptation, transliteration, dialectal translation, pivoting, closely related and morphologically rich languages), eye-tracking for MT evaluation, spoken language translation and speech synthesis. His recent work focuses on analyzing contextualized representations with the focus of linguistic interpretation, manipulation, feature selection and model distillation. His work on analyzing deep neural networks has been published at venues like Computational Linguistics, *ACL, AAAI and ICLR. Nadir has been involved in co-organizing workshops such as simultaneous machine translation and WMT 2019/2020 Machine translation robustness task. He regularly serves as program committee and has served as Area chair at ACL and AAAI this year. 

\bibliography{anthology,eacl2021}
\bibliographystyle{acl_natbib}




\end{document}